OpenStudy (anonymous):

I need someone to walk me through this integral problem:

OpenStudy (anonymous):

\[\int\limits_{0}^{3x} \sqrt{4+t^2 dt}\]

OpenStudy (anonymous):

I am supposed to find f'(x) but I am not sure what to do.

OpenStudy (dumbcow):

since you only need f'(x) the actual integrating is irrelevant assume the antiderivative of sqrt(4+t^2) exists, call it G(t) \[f(x) = |_0^{3x} G(t) = G(3x) - G(0)\] differentiate with respect to x \[f'(x) = g(3x) - 0\] \[f'(x) = \sqrt{4 + (3x)^2}\]

OpenStudy (anonymous):

So I don't have to integrate it, I just have to put 3x and 0 into f'(x)=g(3x)-0

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Do I do that with all of the integral problems that I am taking the derivative of?

OpenStudy (cwrw238):

i was going to integrate it but no need I just been to look up my notes as I had forgotten how to integrate this one

OpenStudy (dumbcow):

oh crap i forgot about the chain rule you also have to multiply by derivative of "3x" \[** f'(x) = 3 \sqrt{4 + (3x)^2}\]

OpenStudy (anonymous):

Is finding the derivative of an integral the same thing as finding that ant-derivative?

OpenStudy (dumbcow):

no the anti derivative is same as the indefinite integral so taking derivative of integral leaves you with original function however when there are limits on the integral it may leave you with a slightly different function based on the limits

OpenStudy (anonymous):

Okay so the derivative of the integral of f sin(x) is just f sin(x).

OpenStudy (anonymous):

The integral is f -cos(x) and so the derivative of that is f sin(x)?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

What is the point to taking the derivative of the integral?

OpenStudy (dumbcow):

however \[\frac{d}{dx} \int\limits_0^{3x} \sin t dt = 3 \sin (3x)\]

OpenStudy (anonymous):

Thank you so much! You have been very helpful!

OpenStudy (dumbcow):

yw glad to help

OpenStudy (anonymous):

I think I will be able to do the next few problems. :)